# Flow-based Conformal Prediction for Multi-dimensional Time Series

### v2

@@ B.1 Experiment Setup
We implemented OT-CP using the source code released by the authors [thurinoptimal]. The training and validation sets were combined to form a calibration set. Following the setup in the original publication, 75% of the calibration set was used to solve OT, and the remaining 25% was used to calibrate the prediction sets.
@@
As the source code from the original publication was not released, we implemented CONTRA ourselves following the methodology and details provided in fang2025contra. Consistent with the original setup, we used six coupling layers with a hidden dimension of 128 and trained the model for 100 epochs with the same batch size as FCP and a learning rate of 0.001. The training and validation sets were combined into a calibration set, of which 50% was used to train the model and the remaining 50% was used to calibrate the prediction sets.
@@
We implemented MultiDimSPCI using the source code released by the authors [xu2024conformal]. The context window size was set to $50$ for all real-world datasets, consistent with the setup used for FCP. Following the original publication, the number of trees was set to $15$. The training and validation sets were combined into a single training set.
@@
We implemented this method using the source code released by the authors [messoudi2021copula], following the setup described in the original publication. The training and validation sets were combined to form a calibration set.
@@
We implemented this method using the source code provided by the authors [messoudi2022ellipsoidal]. Following the setup in the original publication, the training set was used as the proper training set and the validation set as the calibration set. The number of neighbors for kNN was set to 5% of the proper training set size, as suggested by the authors. We also experimented with different neighbor ratios, but these variations did not lead to meaningful differences in performance.
@@
We implemented this method using the source code provided by the authors [sun2022copula], following the setup described in the original publication. The training and validation sets were combined to form a calibration set.
@@
We implemented Temporal Fusion Transformer (TFT) [lim2021temporal] using pytorch_forecasting. A hyperparameter grid search was conducted on the training set of each dataset with $d_{y}=2$ to determine the optimal configuration. We believe this hyperparameter search generalizes well to higher $d_{y}$ within each dataset, since TFT makes predictions for each outcome dimension independently in our setup. Performance was observed to saturate at a model dimension of 32, with two attention heads and two layers, therefore these settings were used for all experiments. For consistency with FCP, the context window size was fixed at 50 across all experiments. We trained the models using the Adam optimizer [kingma2014adam] with a learning rate of 0.001, a maximum of 50 epochs, and a dropout rate of 0.1. Quantile loss with $q\in\{0.025,0.975\}$ was used for 0.95 target coverage.
@@
We implemented DeepAR [salinas2020deepar] using pytorch_forecasting. A hyperparameter grid search was conducted on the training set of each dataset with $d_{y}=2$ to determine the optimal configuration similarly to TFT. Performance was observed to saturate at a model dimension of 32 with two layers, therefore these settings were used for all experiments. For consistency with FCP, the context window size was fixed at 50 across all experiments. We trained the models using the Adam optimizer [kingma2014adam] with a learning rate of 0.001, a maximum of 50 epochs, and a dropout rate of 0.1. Multivariate normal distribution loss with $q\in\{0.025,0.975\}$ was used for 0.95 target coverage.
@@
We used multilayer perceptions (MLP) to model the guided vector field $u_{t|h}:[0,1]\times\mathbb{R}^{d_{h}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}^{d_{y}}$. The time variable $t\in[0,1]$ was concatenated with the input and fed into the vector field. A hyperparameter grid search was conducted on the training set of each dataset with different $d_{y}$ to determine the optimal configuration. We set the hidden dimension of the vector field identical to the model dimension of the encoder, so that additional layer is not required between the vector field and the encoder. Table 3 shows the hyperparameter search space and Table 4 shows the optimized hyperparameter configuration. The context window size for the encoder was set to 50. We trained the model with Adam optimizer [kingma2014adam] with a maximum of 50 epochs for all experiments and used the validation set to select the best model.
To determine an appropriate sample size $N$ for the set size estimation using quasi-Monte Carlo sampling, we computed the relative standard error of the Jacobian determinants of $\psi$, defined as $\textrm{SE}(\det J_{\psi,h})/\textrm{Avg}(\det J_{\psi,h})$, where $\det J_{\psi,h}=\{\det J_{\psi}(x_{j}\mid h)\}_{j=1}^{N}$ are the sampled Jacobian determinants conditioned on $h$. We selected the smallest $N$ such that the average relative standard error across all $h$ falls below 0.01. We used $N=4096$ for experiments with $d_{y}=2$, $N=8192$ for experiments with $d_{y}=4$, and $N=16384$ for experiments with $d_{y}=8$.
Table 4: The optimized hyperparameter configuration for FCP based on the grid search.
@@ B.2 Computational Cost
Table 5 reports the wall-clock training time for all methods, computed as the sum over five independent runs on five different sequences. All models were trained on a machine equipped with dual Intel Xeon Gold 6226 CPUs and a single NVIDIA A100 GPU. For methods that do not employ neural networks, only the CPU was used.
Evaluation time.
Appendix C Dataset Details
Wind dataset
The wind dataset contains wind speed records measured at 30 different wind farms [zhu2021multi]. Each wind farm location provides 768 records with 5 features at each timestamp. We randomly select $d_{y}\in\{2,4,8\}$ locations to construct five sequences of $d_{y}$-dimensional time series.
Traffic dataset
The traffic dataset contains traffic flow collected at 15 different traffic sensor locations [xu2021conformalanomaly]. Each sensor location provides 8778 observations with 5 features at each timestamp. We randomly select $d_{y}\in\{2,4,8\}$ locations to construct five sequences of $d_{y}$-dimensional time series.
Solar dataset
The solar dataset considers solar radiation in Diffused Horizontal Irradiance (DHI) units at 9 different solar sensor locations [zhang2021solar]. Each location provides 8755 records with 5 features at each timestamp. For the solar dataset, we randomly selected $d_{y}\in\{2,4\}$ locations to construct five sequences of $d_{y}$-dimensional time series. We did not construct sequences with $d_{y}=8$ due to the limited number of unique locations, which could lead to overlapping sequences across different trials of experiments.
Table 5: the wall-clock training time (hrs) for all methods.
@@ Appendix C Dataset Details
Table 6: Evaluation time via solving ODE.
Dataset Method $d_{y}{=}2$ $d_{y}{=}4$ $d_{y}{=}8$
Wind FCP $\leq 0.5$ $\leq 0.5$ $\leq 1$
Traffic FCP $\leq 0.5$ $\leq 1.5$ $\leq 5$
Solar FCP $\leq 0.5$ $\leq 1.5$
@@ Appendix D Additional Experiments
D.1 Experiment at 0.9 confidence level
@@
Table 7: Average empirical coverage and prediction sets sizes obtained by FCP and all baselines on three real-world datasets, evaluated under different base predictors and varying outcome dimensions $d_{y}$. Reported values represent the average and standard deviation over five independent experiments. The target confidence level was set to 0.9.
@@
D.2 Rolling Coverage on Wind Dataset
Since conditional coverage is challenging to evaluate in real-world data, we use rolling coverage to approximate conditional coverage at a specific time index. Rolling coverage at time index $i$ is defined as:
@@ D.2 Rolling Coverage on Wind Dataset
where $m$ is a rolling window size. Figure 3 presents the rolling coverage of the test set with rolling window size $m=20$ on the wind dataset.
@@
Figure 3: Rolling coverage results on the wind dataset with rolling window size 20.
D.3 Ablation study
Ablation study on vector field under bi-Lipschitz flow assumption.
Table 8 reports the average empirical coverage and prediction set sizes of FCP with MLP and iResNet across the three datasets with varying $d_{y}$.

### v3

@@ B.1 Experiment Setup
We implement OT-CP using the source code released by the authors [44]. The training and validation sets are combined to form a calibration set. Following the setup in the original publication, 75% of the calibration set is used to solve OT, and the remaining 25% is used to calibrate the prediction sets.
@@
As the source code from the original publication [16] was not released, we implement CONTRA following the methodology and details provided in the original publication. Consistent with the original setup, we use six coupling layers with a hidden dimension of 128 and train the model for 100 epochs with the same batch size as FCP and a learning rate of 0.001. The training and validation sets are combined into a calibration set, of which 50% is used to train the model and the remaining 50% is used to calibrate the prediction sets.
@@
We implement MultiDimSPCI [50] using the source code released by the authors. The context window size is set to $50$ for all real-world datasets, consistent with the setup used for FCP. Following the original publication, the number of trees for the quantile random forest is set to $15$. The training and validation sets are combined into a single training set.
@@
We implement the method using the source code released by the authors [30]. Following the setup described in the original publication, the training and validation sets are combined to form a calibration set.
@@
We implement the method using the source code provided by the authors [31]. Following the setup in the original publication, the training set is used as the proper training set and the validation set as the calibration set. The number of neighbors for kNN is set to 5% of the proper training set size, as suggested in the original publication. We also conduct experiments with different numbers of neighbors, but does not lead to meaningful differences in performance.
@@
We implemented CopulaCPTS using the source code provided by the authors [43]. Following the setup described in the original publication, the training and validation sets are combined to form a calibration set.
@@
We implement Temporal Fusion Transformer (TFT) [28] using Pytorch Forecasting. A hyperparameter grid search is conducted on the training set of each dataset with $d_{y}=2$ to determine the optimal configuration. We believe this hyperparameter search generalizes well to higher $d_{y}$ within each dataset, since TFT makes predictions for each outcome dimension independently in our setup. Performance is observed to saturate at a model dimension of 32, with two attention heads and two layers, therefore this setting is used for all experiments. For consistency with FCP, the context window size is fixed at 50 across all experiments. We train the models using the Adam [25] with a learning rate of 0.001, a maximum of 50 epochs, and a dropout rate of 0.1. Quantile loss with $q\in\{0.025,0.975\}$ is used for 0.95 target coverage.
@@
We implement DeepAR [38] using Pytorch Forecasting. A hyperparameter grid search is conducted on the training set of each dataset with $d_{y}=2$ to determine the optimal configuration similarly to TFT. Performance is observed to saturate at a model dimension of 32 with two layers, therefore this setting is used for all experiments. For consistency with FCP, the context window size is fixed at 50 across all experiments. We train the models using the Adam [25] with a learning rate of 0.001, a maximum of 50 epochs, and a dropout rate of 0.1. Multivariate normal distribution loss with $q\in\{0.025,0.975\}$ is used for 0.95 target coverage.
@@
We use multilayer perceptions (MLP) to model the guided vector field. The time variable $t\in[0,1]$ and guidance $h\in\mathbb{R}^{d_{h}}$ are concatenated with the input and fed into the guided vector field. A hyperparameter grid search is conducted on the training set of each dataset with different $d_{y}$ to determine the optimal configuration. We set the hidden dimension of the vector field identical to the model dimension of the encoder, so that additional layer is not required between the vector field and the encoder. Table 3 shows the hyperparameter search space and Table 4 shows the optimized hyperparameter configuration. The context window size for the encoder is set to 50. We train the model with Adam [25] with a maximum of 50 epochs for all experiments and use the validation set to select the best model.
To determine an appropriate sample size $N$ for the set size estimation using quasi-Monte Carlo sampling, we compute the relative standard error of the Jacobian determinants of $\psi$, defined as $\textrm{SE}(\det J_{\psi,h})/\textrm{Avg}(\det J_{\psi,h})$, where $\det J_{\psi,h}=\{\det J_{\psi}(x_{j}\mid h)\}_{j=1}^{N}$ are the sampled Jacobian determinants conditioned on $h$. We select the smallest $N$ such that the average relative standard error across all $h$ falls below 0.01. As a result, we use $N=4096$ for experiments with $d_{y}=2$, $N=8192$ for experiments with $d_{y}=4$, and $N=16384$ for experiments with $d_{y}=8$.
Table 4: The optimized hyperparameter configuration for FCP based on the hyperparameter search.
@@ B.2 Computational Cost
Table 5 reports the wall-clock training time for all methods, computed as the sum over five independent runs. All models were trained on a machine equipped with dual Intel Xeon Gold 6226 CPUs and a single NVIDIA A100 GPU. For the methods that do not use neural networks, only CPU was used.
Table 5: The wall-clock training time (hrs) for all methods, computed as the sum over five independent runs.
@@
Appendix C Dataset Details
Wind dataset
The wind dataset contains wind speed records measured at 30 different wind farms [56]. Each wind farm location provides 768 records with 5 features at each timestamp. We randomly select $d_{y}\in\{2,4,8\}$ locations to construct five sequences of $d_{y}$-dimensional time series.
Traffic dataset
The traffic dataset contains traffic flow collected at 15 different traffic sensor locations [51]. Each sensor location provides 8778 observations with 5 features at each timestamp. We randomly select $d_{y}\in\{2,4,8\}$ locations to construct five sequences of $d_{y}$-dimensional time series.
Solar dataset
The solar dataset considers solar radiation in Diffused Horizontal Irradiance (DHI) units at 9 different solar sensor locations [55]. Each location provides 8755 records with 5 features at each timestamp. For the solar dataset, we randomly selected $d_{y}\in\{2,4\}$ locations to construct five sequences of $d_{y}$-dimensional time series. We did not construct sequences with $d_{y}=8$ due to the limited number of unique locations, which could lead to overlapping sequences across different trials of experiments.
@@ Appendix D Additional Experiments
D.1 Experiments at Different Confidence Levels
@@
Table 6: Average empirical coverage and prediction sets size obtained by FCP and all baselines on three real-world datasets, evaluated under different base predictors and varying outcome dimensions $d_{y}$. Reported values represent the average and standard deviation over five independent experiments conducted on five constructed sequences. The target confidence level is set to 0.9.
@@
D.2 Rolling Coverage
Since conditional coverage is difficult to assess on real-world data, we use rolling coverage as a proxy for conditional coverage at a specific time index. The rolling coverage at time index $i$ is defined as:
@@ D.2 Rolling Coverage
where $m$ is a rolling window size. Figure 3 presents boxplots of $\widehat{\mathrm{RC}}_{i}$ computed with $m=20$ on the wind dataset.
@@
Figure 3: Boxplots of the rolling coverages on the wind dataset with rolling window size 20 for FCP, OT-CP, and MultiDimSPCI.